# Optimizing Fine-Tuning through Advanced Initialization Strategies for Low-Rank Adaptation

### v1

@@ Abstract.
The rapid development of parameter-efficient fine-tuning methods has noticeably improved the efficiency of adapting large language models. Among these, LoRA has gained widespread popularity due to its strong balance of effectiveness and parameter efficiency. However, LoRA relies on initializing two low-rank matrices whose product is zero, which limits its ability to effectively activate and leverage the original model weights—creating a potential bottleneck for optimal performance. To address this limitation, we propose IniLoRA, a novel initialization strategy that initializes the low-rank matrices to closely approximate the original model weights. Experimental results indicate that IniLoRA achieves better performance than LoRA across a range of models and tasks. Additionally, we introduce two variants, IniLoRA-$\alpha$ and IniLoRA-$\beta$, both leveraging distinct initialization methods to enhance performance further.
@@
$c$s: Computing methodologies Machine learningccs: Machine learning Machine learning approachesccs: Machine learning approaches Neural networks
@@ 1. Introduction
Large language models (LLMs) have advanced natural language processing and multimodal tasks through pre-training on large datasets and fine-tuning with limited labeled data (Zhao et al., 2023; Bang et al., 2023; Liu et al., 2024a; Devlin et al., 2019). While full parameter fine-tuning updates all model parameters for task adaptation, parameter-efficient fine-tuning (PEFT) methods like LoRA reduce memory and compute demands by adjusting only a subset, often through low-rank updates (Hu et al., 2022; Valipour et al., 2023; Pan et al., 2024; Liu et al., 2024b). However, LoRA generally underperforms compared to full fine-tuning, particularly in specialized domains such as coding and complex mathematical reasoning (Biderman et al., 2024).
Figure 1. Our IniLoRA method extends the LoRA framework by introducing an advanced initialization for the low-rank matrices $A$ and $B$.
@@
To investigate the influence of weight initialization on performance, we propose two novel methods. First, IniLoRA-$\alpha$, which employs a broader initialization distribution, is designed to assess the role of initialization in fine-tuning. This variant consistently surpasses both standard LoRA and IniLoRA across benchmarks, highlighting the importance of initialization strategies in fully realizing the potential of parameter-efficient fine-tuning. Second, we introduce IniLoRA-$\beta$, which is based on the Kaiming distribution and further explores the effect of initialization distributions on model performance.
@@ 3. Method
When weight approximation iterations are set to zero, IniLoRA relies on randomly initialized low-rank matrices. Unlike LoRA’s zero-initialized product, this leads to a non-zero product. As shown in Table 7 in Appendix, IniLoRA still achieves competitive results on the GSM8K dataset under this setting with the LLaMA2-7B model, directly challenging LoRA’s initialization strategy. Motivated by this observation, we propose IniLoRA-$\alpha$, which further improves initialization by adjusting the standard deviation. Specifically, IniLoRA-$\alpha$ employs a broader initialization distribution for the low-rank matrices $A$ and $B$, using
$A,B\sim\mathcal{N}(0,\sigma_{\alpha}^{2}),\quad\sigma_{\alpha}>\sigma_{\text{init}}$
@@ 4.1. Weight approximation experiments
To improve the efficiency of weight approximation and optimize computational resource usage, this experiment adopts a concurrent strategy that performs weight approximation across multiple model layers simultaneously. Table 5 in Appendix presents the time and memory requirements of the model evaluated in the experiment. The experiments were conducted on an NVIDIA 4090 GPU with a concurrency level of 64 and a rank value of 8. The results show that applying weight approximation to models with many layers or to self-attention layers with higher hidden dimensions substantially increases both memory consumption and computational time. In contrast, when the model employs a grouped-query attention mechanism, the number of weight parameters in the value module is considerably reduced.
@@ 4.2.3. IniLoRA-$\alpha$ and IniLoRA-$\beta$: Performance on NLG
Model Method Rank 8 Rank 16

### v3

@@ Abstract.
The rapid development of parameter-efficient fine-tuning methods has noticeably improved the efficiency of adapting large language models. Among these, LoRA has gained widespread popularity due to its strong balance of effectiveness and parameter efficiency. However, LoRA relies on initializing two low-rank matrices whose product is zero, which limits its ability to effectively activate and leverage the original model weights—creating a potential bottleneck for optimal performance. To address this limitation, we propose IniLoRA, a novel initialization strategy that initializes the low-rank matrices to closely approximate the original model weights. Experimental results indicate that IniLoRA achieves better performance than LoRA across a range of models and tasks. Additionally, we introduce two variants, IniLoRA-$\alpha$ and IniLoRA-$\beta$, both leveraging distinct initialization methods to enhance performance further. Code is available at https://github.com/hihi885/IniLoRA
@@
$c$s: Computing methodologies Neural networks
@@ 1. Introduction
Parameter-efficient fine-tuning (PEFT) methods, such as LoRA, reduce memory and computational overhead by updating only a subset of model parameters, typically through low-rank adaptations (Hu et al., 2022; Valipour et al., 2023; Pan et al., 2024; Liu et al., 2024). However, research on low-rank matrix initialization remains limited. PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024) explore different strategies for initializing low-rank matrices based on singular value decomposition (SVD). LoftQ (Li et al., 2024) jointly optimizes quantized weights and low-rank matrices to enhance fine-tuning performance in quantized models. Beyond these approaches, we propose several novel strategies for low-rank matrix initialization.
Figure 1. Our IniLoRA method and its two variants extend the LoRA framework by introducing an advanced initialization for the low-rank matrices $A$ and $B$.
@@
To investigate the influence of weight initialization on performance, we propose two novel methods. First, IniLoRA-$\alpha$, which employs a broader initialization distribution, is designed to assess the role of initialization in fine-tuning. This variant consistently surpasses both standard LoRA and IniLoRA across benchmarks, highlighting the importance of initialization strategies in fully realizing the potential of parameter-efficient fine-tuning. Second, we introduce IniLoRA-$\beta$, which is based on the Kaiming distribution (He et al., 2015) and further explores the effect of initialization distributions on model performance.
@@ 3. Method
When weight approximation iterations are set to zero, IniLoRA relies on randomly initialized low-rank matrices. Unlike LoRA’s zero-initialized product, this leads to a non-zero product. As shown in Table 7 in the Appendix, IniLoRA still achieves competitive results on the GSM8K dataset under this setting with the LLaMA2-7B model, directly challenging LoRA’s initialization strategy. Motivated by this observation, we propose IniLoRA-$\alpha$, which further improves initialization by adjusting the standard deviation. Specifically, IniLoRA-$\alpha$ employs a broader initialization distribution for the low-rank matrices $A$ and $B$, using
(4) $A,B\sim\mathcal{N}(0,\sigma_{\alpha}^{2}),\quad\sigma_{\alpha}>\sigma_{\text{init}}$
@@ 4.1. Weight approximation experiments
To improve the efficiency of weight approximation and optimize computational resource usage, this experiment adopts a concurrent strategy that performs weight approximation across multiple model layers simultaneously. Table 5 in the Appendix presents the time and memory requirements of the model evaluated in the experiment. The experiments were conducted on an NVIDIA 4090 GPU with a concurrency level of 64 and a rank value of 8. The results show that applying weight approximation to models with many layers or to self-attention layers with higher hidden dimensions substantially increases both memory consumption and computational time. In contrast, when the model employs a grouped-query attention mechanism, the number of weight parameters in the value module is considerably reduced.
@@ 4.2.3. IniLoRA-$\alpha$ and IniLoRA-$\beta$: Performance on NLG
Model Method Rank 8 Rank 16 Rank 32